OpenStudy (loser66):

Find radius of convergence of \(\sum_{n=0}^\infty \dfrac{5^n}{n!} z^n\) . I got infinitive am I right? Please, help

OpenStudy (thomas5267):

Is this the formula for the radius of convergence? \[ \limsup_{n\to\infty}\sqrt[n]{\frac{5^n}{n!}}z \]

OpenStudy (thomas5267):

Actually no.

OpenStudy (freckles):

what about ratio test

OpenStudy (thomas5267):

I thought he wanted to show that the radius of convergence is infinite. I have not taken any analysis course but Wikipedia says that ratio test only shows that the radius of convergence is finite.

OpenStudy (freckles):

\[n \rightarrow \infty \\ |\frac{a_{n+1}}{a_n}|<1 \\ |\frac{5^{n+1}}{(n+1)!}z^{n+1} \cdot \frac{n!}{5^{n} } \frac{1}{z^n}|<1 \\ |\frac{5}{(n+1)} z|<1 \\ | z| |\frac{5}{n+1}|<1 \\ \text{ remember } n \rightarrow \infty \] maybe you are right

OpenStudy (freckles):

like because there we would get |z|*0<1 which is true 0<1 for all z

OpenStudy (freckles):

oh so nevermind I think it works

OpenStudy (thomas5267):

But it only shows that it is absolutely convergent but not necessarily analytic over the Reals right?

OpenStudy (freckles):

I don't know about all of that. I just know it shows it converges for all z since we have |z|*0<1 is true for all z therefore the radius of convergence is infinite there is a similar example here: http://tutorial.math.lamar.edu/Classes/CalcII/PowerSeries.aspx in example 4....

OpenStudy (freckles):

they do use root test there though instead of ratio test

OpenStudy (loser66):

Unfortunately, my Prof doesn't accept ratio test. We Must do Hamadard test.

OpenStudy (loser66):

\(limsup \sqrt[n]\dfrac{5^n}{n!} = limsup \dfrac{5}{n/e} = limsup \dfrac{5e}{n} = \infty \) Hence \(R = \dfrac {1}{\infty}= ?\) I don't know how to argue then, since it is undefined.

OpenStudy (loser66):

@misty1212

OpenStudy (thomas5267):

\[ \limsup_{n\to\infty} \sqrt[n]{\dfrac{5^n}{n!}} = \limsup_{n\to\infty} \dfrac{5}{n/e} = \limsup_{n\to\infty} \dfrac{5e}{n} = 0 \] Since n is to infinity right?

OpenStudy (thomas5267):

I am not so sure about dropping the \(\sqrt{2\pi n}\) part of the Stirling's approximation.

OpenStudy (loser66):

But then how to argue for R?

OpenStudy (loser66):

R is a number, not limit anymore. 1/0 is undefined

OpenStudy (loser66):

Can we say the radius of convergence is infinitive?

OpenStudy (thomas5267):

If the sequence values are unbounded so that the lim sup is ∞, then the power series does not converge near a, while if the lim sup is 0 then the radius of convergence is ∞, meaning that the series converges on the entire plane. https://en.wikipedia.org/wiki/Cauchy%E2%80%93Hadamard_theorem

OpenStudy (loser66):

Yes, I got you. I know that the radius of convergence is infinitive is different from divergence.

OpenStudy (loser66):

Thank you so much.